# Efficient weighted-ensemble network simulations of the SIS model of epidemics

### v1

@@ III.1 Methodology
Two sources of uncertainty contribute in determining the overall error in the MTE: (i) stochastic nature of the WE method, which produces different results for the same network, and (ii) variability due to different network realizations. To assess the former, we ran WE simulation with a varying number of bins, realizations per bin, and resampling times. The standard deviation of these results served as the network confidence bounds. To assess the latter, we applied the WE method on multiple network realizations. The mean over all networks provided the overall MTE, while the standard deviation of the confidence bounds provided the MTE’s error bars.
@@
In Fig. 2 we compare the runtimes of the KMC and WE methods. Here and in all other figures, the symbol size represents the maximal error bar. One can see that while the KMC scales exponentially with $N$, the WE method is significantly faster, scaling linearly with $N$; e.g., if for $N=10^{3}$ the runtimes were equal, for $N=10^{4}$, the WE method out-competes the KMC by a factor of $\sim 10^{4}$. In Fig. 2 inset we also compare the MTE results of the two methods, confirming the WE method’s accuracy.
@@
Figure 2: A log-log plot of the runtime to obtain the MTE for each of the methods, versus the population size, $N$, normalized by the runtime at $N=10^{3}$, for a gamma-distributed network. Here $k\right>=10$, $\epsilon=3.0$, and $R_{0}=1.3$. The triangles and circles represent the KMC and WE methods, respectively. The KMC scales exponentially with $N$ (blue line – exponential fit), while the WE method scales linearly with $N$ (red line – linear fit). The inset displays the logarithm of the MTE as a function of $N$, with circles and triangles representing the results of the KMC and WE methods, respectively, for the same parameters as in the main figure. Circles denote WE simulations results, while triangles denote KMC results.
III.2 Results
@@
We employed the WE method to compute the MTE for four different networks, each with a different degree distribution. Results for the gamma distribution are shown in Fig. 2, while results for the inverse-Gaussian, beta and log-normal distributions as a function of population size are shown in Fig. 3. Since we expect the MTE to have the following scaling: $T_{exp}\simeq A\,N^{\alpha}\,e^{N{\cal S}}$ [19, 27], we have fitted the logarithm of the MTE to a linear function of $N$ with logarithmic corrections. First of all, the WE method allows simulating huge networks of size $N=10^{5}$ and more, which is well beyond the capabilities of the KMC method. Second, the scaling of the MTE with $N$ holds for heterogeneous networks as well – here we have taken the COV $\epsilon$ to be very high, and still the simple formula, which was developed for the well-mixed case, qualitatively holds. Third, even though there is an offset between networks, one can see that the logarithmic slopes are similar across networks with similar degree distributions: identical mean and COV.
@@
Figure 3: Log-Log plot of the logarithm of MTE versus the population size, $N$, for networks with inverse-Gaussian (triangles), beta (squares) and log-normal (diamonds) degree distributions. Here, $k\right>=10$, $\epsilon=3.0$, and $R_{0}=1.3$. Solid lines depict linear fits with logarithmic corrections, see text.
@@
The impact of the coefficient of variation on the MTE is shown in Fig. 4, where the MTE is plotted against the square of the coefficient of variation. Notably, and as also seen in Fig. 3, the WE method allows us to obtain very long extinction times (here, up to $10^{130}$), well beyond the capabilities of the KMC method. Our results indicate that the MTE strongly decreases with increasing $\epsilon$; namely, as heterogeneity increases, disease clearance becomes more and more likely, even for huge networks. For $\epsilon^{2}=8$ (highly skewed networks, see inset), the MTE drops dramatically. One can conjecture that the action barrier multiplying $N$ in the exponent, see Eq. (5), vanishes as $\epsilon$ increases. This can be explained in a straightforward manner: as the COV increases, there are more and more high-degree hubs. Once these recover, their neighbors’ infection rate decreases which effectively decreases $R_{0}$, making extinction more likely.
@@
For weak heterogeneity we also compared our numerical results in Fig. 4 to previous analytical results obtained for heterogeneous networks with weak heterogeneity, see Eq. (II[25]. Notably, the WE method allows to simulate huge networks with varying heterogeneity, and allows to plot the dependence of the MTE on $\epsilon$ for a wide range of COVs, which was inaccessible using the KMC method. Interestingly, given a network size, when heterogeneity is weak, the results for the MTE depend only on the mean and COV of the degree distribution, and are independent on its higher moments, as predicted by [25]. However, as $\epsilon$ increases, the different network results depart from each other, as they are no longer independent on the higher moments of the degree distribution, see inset of Fig. 4.
@@
Figure 4: The MTE versus COV squared, $\epsilon^{2}$, for networks with gamma (circles), inverse-Gaussian (triangles), beta (squares) and log-normal (diamonds) degree distributions. Here, $N=10^{4}$, $k\right>=20$, and $R_{0}=1.3$. The star denotes the MTE for a network with exponential degree distribution, with $\epsilon=1$, while the solid line represents the analytical prediction for weak heterogeneity [Eq. (II)]. Inset shows the large-$\epsilon$ regime where the results for the various networks depart.
@@
While the network size and COV play a crucial role in determining the MTE, other parameters may also be key in determining the extinction dynamics. Such is the basic reproduction number $R_{0}$. In Fig. 5, the natural logarithm of the MTE is plotted as a function of $R_{0}$ for networks with different degree distributions but with the same mean degree and COV. First of all, as expected, the figure shows a dramatic increase in the MTE as $R_{0}$ increases, with a huge variability across networks for larger values of $R_{0}$. Since we demand that the first two moments of the networks be identical, the variability here probably originates from the difference in the skewness of the networks, which strongly impacts rare events. The inset of Fig. 5 shows the MTEs close to bifurcation, where $R_{0}-1\ll 1$. Here we find a good agreement between our numerical results and results found for arbitrary heterogeneous networks close to bifurcation, see Eq. (6[27]. Notably, the significant differences across networks with the same mean and COV, emphasize the need for reliable and fast numerical schemes, such as the WE method, to account for the complex interplay between topological and dynamical features.
IV Discussion
We have studied the long-time dynamics and disease extinction in the realm of the SIS model of epidemics using the so-called weighted-ensemble (WE) method. In most previous works, the SIS dynamics on population networks has been analyzed using the “brute-force” Gillespie kinetic Monte-Carlo (KMC) algorithm. This algorithm is excellent, when one is interested in simulating short time scales related to the emergence of the endemic state and its relaxation dynamics, or to compute the extinction dynamics on small networks. However, the algorithm becomes infeasible computationally for large heterogeneous networks. Naturally, the sequential nature of the KMC prevents the study of rare events such as disease extinction, as the computational time required grows exponentially with the system size. Furthermore, when heterogeneity increases, the huge variability between node degrees makes standard KMC algorithms less efficient, as time steps may become vanishingly small.
@@ IV Discussion
Figure 5: The MTE versus the basic reproduction number, $R_{0}$, for networks with gamma (circles), inverse-Gaussian (triangles), beta (squares) and log-normal (diamonds) degree distributions. Here $N=10^{4}$, $k\right>=20$, and $\epsilon=1.5$. Inset shows a log-log plot of $\ln T_{ext}$ versus $R_{0}-1$ for gamma- (circles) and inverse-Gaussian-distributed (triangles) networks. Here, the solid and dashed lines depict the analytical prediction, Eq. (6), for these networks.
@@
The WE method, which we have introduced here to study rare events on population networks, efficiently explores reaction pathways by generating multiple replicas. As the simulation progresses, replicas that are close to extinction are allocated more computational resources, while others are discarded, thereby enabling us to study large heterogeneous population networks that were previously inaccessible with KMC methods. After demonstrating an equivalence between the KMC and WE methods and showing that the runtime of the latter scales linearly with the system’s size, we studied the dependence of the mean time to extinction (MTE) on three main features: the network’s size $N$, its coefficient of variation $\epsilon$, and the basic reproduction number $R_{0}$. We have found that the dependence on $N$ in strongly heterogeneous networks is qualitatively similar to that of well-mixed systems. We have also found that increasing heterogeneity drastically decreases the disease lifetime. Finally, we have shown that increasing $R_{0}$ causes a surge in the MTE and span across networks with similar characteristics. In addition, we have managed to asymptotically corroborate our results against known analytical results in the limits of weak heterogeneity and close to bifurcation.
@@
While we have focused on random networks, with no degree correlations between nearest neighbors, the WE method can also be used to study assortative networks. These are characterized by the tendency of high-degree nodes to be connected to other high-degree nodes, and vice versa; as a result they tend to percolate more easily, forming a giant component that contains a significant fraction of the entire network’s nodes [51, 52]. It has been recently shown that correlation between neighboring nodes can alter the outbreak threshold [53, 46] and change the outbreak size distribution’s characteristics [54]. Yet, the study of rare events on assortative networks requires large populations sizes, to overcome finite size effects. Thus, the WE method implemented here is ideal for studying the interplay between network heterogeneity and assortativity. Since the latter is prevalent in many types of networks such as social, technological, and biological networks [51, 52, 8], the WE method can play a key role in future studies of complex networks.
V Acknowledgments
@@
V.1 Gamma-distribution
The PDF of a gamma distribution with a shape parameter $\kappa>0$ and scale parameter $\theta>0$ is given by,
@@ V.1 Gamma-distribution
$\theta=\epsilon^{2}\left<k\right>,\quad\kappa=\frac{1}{\epsilon^{2}}\;\;\Longleftrightarrow\;\;\langle k\rangle=\kappa\theta,\quad\epsilon=\frac{1}{\sqrt{\kappa}}.$ (A2)
@@
V.2 Beta-distribution
The beta distribution’s PDF is represented as follows:
$p(x;\alpha,\beta)=\frac{1}{B(\alpha,\beta)}x^{\alpha-1}(1-x)^{\beta-1},\quad 0\leq x\leq 1$ (A3)
@@ V.2 Beta-distribution
Here, to obtain a fat-tailed distribution, one must have $\alpha\lesssim 1$. If one is interested in having both large mean and large COV, one has to choose $\beta/N\ll\alpha\lesssim 1$, which yields distributions such as shown in blue in Fig. S1.
@@
V.3 Wald distribution
The inverse Gaussian distribution, also known as the Wald distribution, has a PDF given by:
@@ V.3 Wald distribution
V.4 Log-normal distribution
Lastly, the log-normal distribution PDF reads,
$p(x;\mu,\sigma)=\frac{1}{x\sigma\sqrt{2\pi}}\exp\left[-\frac{\left(\ln x\!-\!\mu\right)^{2}}{2\sigma^{2}}\right],\quad x>0,$ (A7)
@@ V.4 Log-normal distribution
$\mu$ $=$ $-\frac{1}{2}\ln\left(\frac{1+\epsilon^{2}}{\left<k\right>^{2}}\right),\quad\sigma^{2}=\ln(1+\epsilon^{2})\;\;\Longleftrightarrow$
@@
Here, to obtain a fat-tailed distribution with a large mean, one must have $\sigma\gtrsim 1$ and a large enough $\mu$, such that $\mu+\sigma^{2}/2\gg 1$, which yields distributions such as shown in blue in Fig. S1.

### v2

@@ III NUMERICAL SIMULATIONS
To estimate the error of the WE calculations of the MTE and QSD, we identify two sources of uncertainty: (i) stochastic nature of the method, producing different results for the same network, and (ii) variability due to different network realizations. To assess (i) we ran WE simulations with a varying number of realizations per bin $5000400\!<\!m\!<\!$ 5000, resampling time $20.2\!<\!\tau_{WE}\!<\!$ 2, and a fixed number of time steps $M=70$. The standard deviation of these results served as the network confidence bounds. To assess (ii) we applied the WE method on multiple network realizations. The mean over all networks provided the overall MTE, while the standard deviation of the confidence bounds gave the MTE’s error bars. The latter are captured by the symbol sizes in all figures.
@@
In Fig. 4 we compare the runtimes of the KMC and WE methods. One can see that while the KMC scales exponentially with $N$, the WE method is significantly faster, scaling linearly with $N$; e.g., if for $N=10^{3}$ the runtimes were equal, for $N=10^{4}$, the WE method out-competes the KMC by a factor of $\sim 10^{4}$. The inset of Fig. 4 compares the MTE results of the two methods, confirming the WE method’s accuracy.
@@
Figure 4: A log-log plot of the runtime to obtain the MTE for each of the methods, versus the population size, $N$, normalized by the runtime at $N=10^{3}$, for a gamma-distributed network. Here $k\right>=10$, $\epsilon=3.0$, and $R_{0}=1.3$. The WE parameters are $m=10^{3}$, $M=70$ and $\tau_{WE}=1$ for $N<10^{5}$, while $\tau_{WE}=0.5$ for $N=10^{5}$. Triangles and circles represent the KMC and WE methods, respectively. The KMC scales exponentially with $N$ (blue line – exponential fit), while the WE method scales linearly with $N$ (red line – linear fit). Inset shows the logarithm of the MTE versus $N$. Symbols and parameters are the same as in the main figure.
IV Results
@@
We employ the WE method to compute the MTE for four different networks, each with a different degree distribution. Results for the gamma distribution are shown in Fig. 4, while results for the inverse-Gaussian, beta and log-normal distributions as a function of the population size are shown in Fig. 5. Based on previous work, the general form of the MTE is expected to have the following scaling with $N$: $T_{\text{ext}}\simeq A\,N^{\alpha}\,e^{N{\cal S}}$ [19, 27], where $\alpha$ and $A$ are some constants, and ${\cal S}$ is the action barrier. To corroborate this scaling, we have fitted the logarithm of the MTE to a linear function of $N$ with logarithmic corrections. Notably, the WE method allows simulating very large networks of size $N=10^{5}$ and higher, which is well beyond the capabilities of the KMC method, whose runtime diverges at such network sizes (see Fig. 4). In Fig. 5 we demonstrate that this scaling of the MTE with $N$, derived for homogeneous and weakly-heterogeneous network, also holds for strongly heterogeneous networks, with a very high COV. Even though there is an offset between different networks (indicating a different constant $A$), as long as the mean and COV are fixed, one observes good qualitative agreement across different networks.
@@
Figure 5: Log-Log plot of the logarithm of MTE versus $N$, for inverse-Gaussian (triangles), beta (squares) and log-normal (diamonds) networks. Here, $k\right>=10$, $\epsilon=3.0$, $R_{0}=1.3$, and the WE parameters are $M=70$, $m=500$ and $\tau_{WE}=1$ for $N<10^{5}$, and $m=10^{3}$ and $\tau_{WE}=0.5$ for $N=10^{5}$. Solid lines are linear fits with logarithmic corrections, see text.
@@
The impact of the COV, $\epsilon$, on the MTE is shown in Fig. 6, where the MTE is plotted against $\epsilon^{2}$. Notably, the WE method allows us to simulate the dynamics on large networks over a large span of COVs, which yield a huge range of MTEs, well beyond the capabilities of the KMC method. Our results in Fig. 6 indicate that the MTE strongly decreases with increasing $\epsilon$; namely, as heterogeneity increases, disease clearance becomes more and more likely, even for very large networks. For very strong heterogeneity ($\epsilon=\sqrt{8}$, see inset), the MTE drops drastically. Here, apparently the action barrier multiplying $N$ in the exponent, see Eq. (5), vanishes as $\epsilon$ increases. The reason for this is straightforward: as the COV increases, there are more and more high-degree hubs. Once these recover, their neighbors’ infection rate decreases which effectively decreases $R_{0}$, making extinction more likely.
@@
For weak heterogeneity we also compared our numerical results in Fig. 6 to Eq. (II), obtained for weakly-heterogeneous networks [25]. Notably, given a weakly-heterogeneous network of size $N$, the MTE is solely determined by the mean and COV of the degree distribution $p(k)$, as the results of all networks coincide there [25]. Yet, as $\epsilon$ increases, the MTEs of different networks depart from each other, as they are no longer independent on the higher moments of $p(k)$, see inset of Fig. 6.
@@
Figure 6: The MTE versus $\epsilon^{2}$, for gamma (circles), inverse-Gaussian (triangles), beta (squares) and log-normal (diamonds) networks. Here, $N=10^{4}$, $k\right>=20$, $R_{0}=1.3$, and WE parameters are $M\!=\!70$, $m\!=\!10^{3}$ and $\tau_{WE}\!=\!1.0$ for simulations with $\epsilon\!>\!0.5$, while $30001000\!<\!m\!<\!$ 3000 and $0.60.2\!<\!\tau_{WE}\!<\!$ 0.6 for simulations with $\epsilon\!<\!0.5$. The star denotes the MTE for a network with exponential degree distribution, with $\epsilon=1$, while the solid line represents the analytical prediction for weak heterogeneity [Eq. (II)]. Inset shows the large-$\epsilon$ regime where the results for the various networks depart.
@@
While the network size and COV play a crucial role in determining the MTE, other parameters may also be key in determining the extinction dynamics. Such is the basic reproduction number $R_{0}$. In Fig. 7, the natural logarithm of the MTE is plotted as a function of $R_{0}$ for different networks with a fixed mean degree and COV. As expected, the figure shows a dramatic increase in the MTE as $R_{0}$ increases, with a huge variability across networks for larger values of $R_{0}$. The inset of Fig. 7 shows the MTEs close to bifurcation, $R_{0}-1\ll 1$, where our results agree well with Eq. (6), obtained for arbitrary heterogeneous networks close to bifurcation [27].
@@
Figure 7: The MTE versus $R_{0}$ for gamma (circles), inverse-Gaussian (triangles), beta (squares) and log-normal (diamonds) networks. Here $N=10^{4}$, $k\right>=20$, $\epsilon=1.5$, and WE parameters $m=10^{3}$, $\tau_{WE}=1$ and $M=70$. Inset shows a log-log plot of $\ln T_{\text{ext}}$ versus $R_{0}-1$ for gamma (circles) and inverse-Gaussian (triangles) networks. The solid and dashed lines depict the analytical prediction (6) for these networks.
Importantly, in Fig. 7, despite the fact that we fixed the network’s first two moments and size, varying $R_{0}$ results in significant variability in the MTE for different network distributions. The same phenomenon occurs in Fig. 6, see inset. The reason is that when the COV is large, $\epsilon\gtrsim 1$, the degree distribution is highly skewed and thus, higher moments become important in determining the MTE. To test this, we plot in Fig. 8 the MTE versus the skewness $\tilde{\mu}_{3}$—the third standardized moment of the distribution. Here, for simplicity, the degree distribution is given by a mixture of two Gaussian distributions with different means and equal variance, allowing us to adjust the distribution’s skewness by controlling the relative heights and spread of each peak. As the skewness $\tilde{\mu}_{3}$ increases from a large negative value to a large positive value, the median degree moves to the left, i.e., becomes smaller. As shown in Fig. 8, the MTE decreases with increasing $\tilde{\mu}_{3}$, despite the fact the distribution’s mean is fixed. This indicates that the MTE decreases with decreasing median, which is a measure for the degree distribution’s typical degree. To further explore this dependence, in the inset of Fig. 8 we compare the MTEs of a calculation with varying median and fixed mean, to a calculation with varying mean and fixed median. One can clearly see that the dependence of the MTE on the median is much more pronounced than on the mean.
The reason for the strong dependence of the MTE on the typical degree is as follows. In heterogeneous networks, the effective reproductive number is largely determined by the hubs, which tend to get infected and also recover quite rapidly. Once a hub recovers, the infection rate of its neighbors drops, thereby effectively decreasing $R_{0}$ as well as the overall MTE. As the skewness increases while keeping the mean and COV fixed, the number of hubs increase while the typical degree decreases. Therefore, a decrease in the median here is a precursor for an increase in the number of hubs, and decrease in the MTE.
Figure 8: MTE normalized by its maximal value, $T_{\text{ext}}^{(0)}$, versus skewness $\tilde{\mu}_{3}$ for a network with $N=10^{4}$ nodes, with degree distribution given by a mixture of two Gaussians, see text. Here $R_{0}=1.2$ and we fix the mean $\langle k\rangle=50$ and the COV $\epsilon=0.6$ (red triangles) and $\epsilon=0.5$ (black triangles). The WE parameters are $m=500$, $\tau_{WE}=0.3$, and $M=70$. Inset shows the MTE normalized by its maximal value versus $\tilde{\mu}_{3}$ for $N=10^{4}$, $R_{0}=1.2$, and $\epsilon=0.6$ using WE parameters $m=500$, $\tau_{WE}=0.5$ and $M=70$. Simulations with fixed mean $\langle k\rangle=50$ and varying median (triangles), are compared with simulations with a fixed median of $50$ and varying $\langle k\rangle$ (circles). All dashed lines represent a linear fit.
V Discussion
We have studied the long-time dynamics and disease extinction in the realm of the SIS model of epidemics using the so-called weighted-ensemble (WE) method. In most previous works, the SIS dynamics on population networks has been analyzed using the “brute-force” Gillespie kinetic Monte-Carlo (KMC) algorithm. This algorithm is highly efficient at short time scales, allowing to study, e.g., the epidemic threshold or stability of the endemic state on very large population networks. However, when one is interested to compute long-time dynamics such as disease clearance on large networks, the KMC algorithm becomes infeasible, with a computation time growing exponentially with the network size. Furthermore, as heterogeneity increases, the large variability between node degrees makes standard KMC algorithms less efficient, as time steps may become vanishingly small.
These reasons emphasize the need for reliable and fast numerical schemes, such as the WE method which we have introduced here, to efficiently study rare events on heterogeneous population networks. In the WE method, as the simulation progresses, replicas that are closer to extinction are allocated more computational resources, while others are discarded, thereby enabling us to accurately compute probabilities that were previously beyond reach with KMC methods. Thus, the WE method allows, e.g., computing the quasi-stationary distribution in the long-lived endemic state including the exponentially small tails, and the mean time to extinction (MTE), by measuring the total flux into the extinction state.
@@ V Discussion
After demonstrating an equivalence between the KMC and WE methods and showing that the runtime of the latter scales linearly with the system’s size, we studied the dependence of the MTE on three main features: the network’s size $N$, its coefficient of variation (COV) $\epsilon$, and the basic reproduction number $R_{0}$. We have found that the dependence on $N$ in strongly heterogeneous networks is qualitatively similar to that of well-mixed systems. We have also shown that increasing heterogeneity drastically decreases the disease lifetime, whereas increasing $R_{0}$ strongly increases the MTE. An important feature that we have revealed is that even if the network’s mean and COV are fixed, there can still be substantial variability in the MTE across different networks. To explain this, we analyzed the impact of degree skewness on the MTE and found that higher skewness leads to shorter disease lifetimes, suggesting that the median (or typical) degree serves as a more reliable indicator of MTE variation than the average degree. Furthermore, we have managed to asymptotically corroborate previous theoretical results, which were obtained for weakly-heterogeneous networks, and when the dynamics is close to bifurcation, in parameter regimes previously inaccessible to simulations.
@@
While we have focused on random networks, with negligible degree-degree correlations between nearest neighbors, the WE method can also be used to study assortative networks. These are characterized by the tendency of high-degree nodes to be connected to other high-degree nodes, and vice versa; as a result they tend to percolate more easily, forming a giant component that contains a significant fraction of the entire network’s nodes [49, 55]. It has been recently shown that correlations between neighboring nodes affect the final outbreak size [56] and also the epidemic threshold; the latter, in general, is proportional to the largest eigenvalue of the adjacency matrix [8, 57, 32, 48]. Yet, the study of rare events on assortative networks requires large network sizes, to overcome finite-size effects. Thus, even if the epidemic threshold can only be found numerically, the WE method is ideal for studying the interplay between network heterogeneity and assortativity. Since assortativity is key in many real-life social, technological, and biological networks [49, 55, 8], the WE method can thus play a crucial role in the analysis of such complex, realistic networks.
VI Acknowledgments
@@
VI.0.1 Gamma-distribution
The gamma PDF with a shape parameter $\kappa>0$ and scale parameter $\theta>0$ is given by:
@@ VI.0.1 Gamma-distribution
$\theta=\epsilon^{2}\left<k\right>,\quad\kappa=1/\epsilon^{2}\;\Longleftrightarrow\;\langle k\rangle=\kappa\theta,\quad\epsilon=\kappa^{-1/2}\!.$ (A2)
@@
VI.0.2 Beta-distribution
The beta PDF is given by:
$p(x;\alpha,\beta)=B(\alpha,\beta)^{-1}x^{\alpha-1}(1\!-\!x)^{\beta-1}\!,\quad 0\leq x\leq 1.$ (A3)
@@ VI.0.2 Beta-distribution
In this case, to get a fat-tailed distribution, one must have $\alpha\lesssim 1$. If one is interested in having both large mean and large COV, one has to choose $\beta/N\ll\alpha\lesssim 1$, which yields distributions such as shown in blue in Fig. S1.
@@
VI.0.3 Wald distribution
The inverse Gaussian (or Wald) PDF reads:
@@ VI.0.3 Wald distribution
VI.0.4 Log-normal distribution
Lastly, the log-normal PDF reads:
$p(x;\mu,\sigma)=\left(x\sigma\sqrt{2\pi}\right)^{\!-\!1}\exp\left[\!-\!\left(\ln x\!-\!\mu\right)^{2}\!/(2\sigma^{2})\right]\!,\;\;x\!>\!0,$ (A7)
@@ VI.0.4 Log-normal distribution
$\mu$ $=$ $-(1/2)\ln\left[(1+\epsilon^{2})/\left<k\right>^{2}\right],\quad\sigma^{2}=\ln(1+\epsilon^{2})\;\;\Longleftrightarrow$
@@
Here, to obtain a fat-tailed distribution with a large mean, one must have $\sigma\gtrsim 1$ and a large enough $\mu$, such that $\mu+\sigma^{2}/2\gg 1$, see distributions in blue in Fig. S1.